If your eye glasses have focal length 60cm what is your near point?
1
Expert's answer
2019-09-06T09:46:19-0400

dn=0.25m

"D_n=1\/0.25=4D"

"D_{glasses}=1\/0.6=1.67D"

"\Delta D=4-1.67=2.33D"

"d=\frac{1}{\Delta D}=\frac{1}{2.33}=0.43m"

Answer d=43cm